# Autoassociative Learning of Structural Representations for Modeling and Classification in Medical Imaging

### v2

@@ 4.4 Stage 1: Image reconstruction
Table 1 compares the performance of configurations in terms of four reconstruction metrics comparing the produced output image to the input image, all calculated on the test set: Mean Absolute Error (MAE), Mean Square Error (MSE), Structural Similarity Index Measure (SSIM) by Wang et al. (2004), and Masked MSE (MMSE), i.e. the actual reconstruction loss used in training (Sec.2.2). For each configuration, five models have been trained, each starting from a different initialization.
Table 1: The performance of individual runs of three variants of ASR and the Baseline autoencoder. Gray background marks the model with the best MSE within a configuration.
MODEL MAE ↓ MSE ↓ SSIM ↑ MMSE ↓
Base_1 0.1393 0.0342 0.1936 0.0342
Base_2 0.1370 0.0330 0.1943 0.0330
Base_3 0.1362 0.0327 0.1938 0.0326
Base_4 0.1367 0.0329 0.1924 0.0329
Base_5 0.1363 0.0326 0.1921 0.0325
Mean 0.1371 0.0331 0.1932 0.0330
Regularized_1 0.1415 0.0341 0.1968 0.0341
Regularized_2 0.1417 0.0341 0.1976 0.0341
Regularized_3 0.1431 0.0350 0.1972 0.0350
Regularized_4 0.1407 0.0342 0.1976 0.0342
Regularized_5 0.1412 0.0345 0.1961 0.0345
Mean 0.1416 0.0344 0.1971 0.0344
Incremental_1 0.1415 0.0342 0.1956 0.0342
Incremental_2 0.1424 0.0347 0.1958 0.0347
Incremental_3 0.1428 0.0349 0.1968 0.0349
Incremental_4 0.1466 0.0366 0.1961 0.0366
Incremental_5 0.1420 0.0346 0.1960 0.0346
Mean 0.1431 0.0350 0.1961 0.0350
Baseline_1 0.1178 0.0274 0.2299 -
Baseline_2 0.1187 0.0276 0.2271 -
Baseline_3 0.1179 0.0272 0.2295 -
Baseline_4 0.1174 0.0271 0.2292 -
Baseline_5 0.1202 0.0278 0.2227 -
Mean 0.1184 0.0274 0.2277 -
@@
Figure 5: Reconstructions of sample test images produced by the REG_2 model.
@@ 4.5 Stage 2: Classification
The encoders produce image features that are subsequently used to train and test a classifier. For ASR, this process comprises the following steps for a given bag $B$:
  1. 1.
    Querying the Encoder and the Modelers on each of the sixteen 256$\times$256 image patches from $B$, each resulting with vectors of latent features per scale and location in grids.
  2. 2.
    Aggregation of the latent features simultaneously over space (the X and Y dimensions of the spatial latent) and over $B$, for each scale (0, 1, 2) and each latent variable independently. For instance, one of such variables is the rotation of the ellipse at scale 0. As there are 6 ellipse-controlling variables at each scale, there are 18 variables in total. These are then aggregated independently using mean and standard deviation, resulting in 36 features.
The result of the above process is a tabular dataset of 1,463 bags (examples, rows) described by 36 attributes (columns) and partitioned into the three considered decision classes. The partitioning of that dataset into training, validation and test subsets is the same as in Sec. 4.1.
@@
The induction algorithm. To demonstrate the explanatory potential of ASR, we choose decision trees as the predictive model to be trained on the tabular datasets prepared in the way described above. We use the DecisionTreeClassifier implementation from the scikit-learn library (Pedregosa et al., 2011) that relies on an optimized version of the CART algorithm (Breiman et al., 1984). In training, we optimize the tree induction algorithm by exhaustively iterating over 30 following combinations of hyperparameter values: node impurity measure (gini, entropy), maximum height of the tree (3, 4, 5, 7, $\infty$), and the minimum number of samples per leaf (5, 10, 20). For each of those combinations, the accuracy of classification is assessed using a 5-fold cross validation conducted on the training part of the tabular dataset. The combination that maximizes this measure is subsequently used to build a decision tree from the entire training set; however, to further maximize the predictive capability, this tree undergoes adaptive pruning using the minimal cost-complexity method. The intensity of pruning is controlled by the ccp_alpha parameter of DecisionTreeClassifier, with 0 being a neutral value (no pruning) and larger values causing folding of terminal subtrees into leaves. We tune ccp_alpha to maximize the classification accuracy on the validation set. The tree pruned in this way is subsequently tested on the test set.
@@
Results. We produce a decision tree in the above way for each ASR and Baseline model listed in Table 1, and asses its predictive capacity on the testing part of the tabular dataset. Table 2 presents the accuracy of classification, precision, recall and F1 score of each of those classifiers. Notably, all configurations of ASR outperform the Baseline models on all metrics, and do so by a large margin. The best Baseline model is not better than the worst ASR model on any metric. One may argue that this may be due the Baseline models having over 5 times more attributes at their disposal (200 vs. 36). However, the extensive hyperparameter tuning provided fair footing for the Baseline configurations, allowing them to pick the optimal hyperparameter values and pruning intensity. Apparently, however, this was not sufficient for Baseline to perform well: the ‘anonymous’ features of its latent representation were not informative enough to provide for good predictive capacity. ASR, in contrast, managed to provide the decision trees with useful diagnostic information – even though at the time of Stage 1 training (Sec. 4.4) it was oblivious about decision classes.
@@
Table 2: Test-set predictive performance of each model. Gray shading marks the best model in terms of both accuracy and F1-score within a given group.
MODEL accuracy ↑ precision ↑ recall ↑ f1_score ↑
Base_1 0.5864 0.6245 0.5864 0.5048
Base_2 0.7768 0.8096 0.7768 0.7735
Base_3 0.7112 0.7164 0.7112 0.7038
Base_4 0.5930 0.5907 0.5930 0.5523
Base_5 0.7615 0.7957 0.7615 0.7414
Regularized_1 0.6980 0.7044 0.6980 0.6738
Regularized_2 0.7133 0.7040 0.7133 0.7040
Regularized_3 0.6827 0.6762 0.6827 0.6655
Regularized_4 0.6674 0.7079 0.6674 0.6785
Regularized_5 0.7133 0.7245 0.7133 0.7075
Incremental_1 0.7090 0.7471 0.7090 0.6892
Incremental_2 0.6958 0.7058 0.6958 0.6998
Incremental_3 0.6521 0.7061 0.6521 0.6131
Incremental_4 0.7505 0.7568 0.7505 0.7511
Incremental_5 0.7396 0.7651 0.7396 0.7339
Baseline_1 0.4486 0.2979 0.4486 0.3572
Baseline_2 0.4311 0.3945 0.4311 0.3797
Baseline_3 0.2801 0.3729 0.2801 0.2342
Baseline_4 0.3129 0.3928 0.3129 0.2776
Baseline_5 0.5383 0.4108 0.5383 0.4486
@@
Statistical analysis of the results presented in Table 2 ANOVA and post-hoc Holm test indicate statistically significant superiority of each variant of ASR compared to the Baseline, on accuracy of classification and on the F1 score (Sec. C). There are no significant differences between the variants of ASR. Notably, the Base variant exhibits most variance between runs, while the Regularized one is most stable in this respect.
@@
Figure 7 presents the decision tree that achieved the best accuracy of classification across all ASR runs (the second run of the Base configuration). Similarly to other trees induced from the ASR-generated latent representation (not shown here for brevity), this tree nicely delineates most examples of the Hashimoto decision class (H) already at the root of the tree, using the mean of the height ($h$) ellipse parameter at scale $j=1$. The remaining two decision classes (N and B) turn out to be much harder to differentiate. Overall, the Nodularity class turns out to be hardest to classify, which we attribute to the relatively large fraction of connective tissue and sparse, small blobs, which are difficult to reconstruct using only ellipses (Fig. 3). Additionally, the examinations from this decision class occasionally feature inflammation, which can be mismatched with some visual features of Hashimoto.
@@
More importantly, the obtained decision tree is notably small, comprising only 6 decision nodes and 7 leaves [^3] It is easy to notice that the pruning algorithm failed to fold the right-hand Hashimoto subtree, which would further reduce these numbers. Other trees induced from attributes provided by ASR were similarly compact. This corroborates the main hypothesis of this study, i.e. that the structural, compositional representation enforced by ASR correlates better with relevant visual features, and provides a solid basis for successful classification. Interestingly, out of the five unique attributes used by decision nodes, only one (the one used by the root node) comes from scale $j=1$, and none come from the scale $j=2$. This signals that the features of large, coarse ellipses built at the lowest scale are most useful for this particular diagnostic scenario. To verify this supposition, we performed feature importance analysis by assessing the gain of the Gini criterion for each attribute (a.k.a. Gini importance), and averaged it over the final, optimized and pruned decision trees obtained in all 15 runs of ASR variants. Figure 8 shows the result of this analysis, which confirms that scale 0 features are most important, followed by scale 1 and then scale 2. In the group of color features (ar, ag, ab), the standard deviation of the green color component at scale 0 turns out to be particularly informative; this may be caused by the transmissive nature of the imaging process, in which changes in the green component result in different shades of purple. Among the geometric features, the standard deviation of the orientation angle of the ellipse (d) at the same scale proves most informative.
Figure 8: The importance of each of 36 attributes averaged over 15 decision trees obtained with ASR.
The above description demonstrates also the transparency of ASR, where decision-making process can be traced back to interpretable attributes expressed in terms of visual concepts that are intuitive for humans. Prospectively, this capability can be further exploited by, for instance, pointing to the geometric primitives (here: ellipses) that contributed most to particular stages of the decision-making process, represented by the individual nodes of the decision tree.
@@ 5 Conclusions
In this study, we provided preliminary evidence for the usefulness of a new representation learning approach, in which the learner is forced to ‘explain’ the data in terms of higher-level concepts available in the rendering process. This allows forming compact, informative and explainable representations, which are not only capable of conveying the essence of the content of the input image, but also constructing potent predictors with high explanatory power.
ASR offers natural means for providing a learning system with domain knowledge in the form of the repertoire of visual primitives. For the class of images considered in this study, elliptical shapes proved sufficient to outperform a conventional DL baseline. It is likely, however, that more sophisticated graphical representations could further improve the quality of reconstruction and the accuracy of the resulting classifiers; for instance, the Fourier-based shape representations used in (Krawiec and Nowinowski, 2024) offer more flexibility in shape modeling, and thus could likely convey more relevant information. For this and other domains, whether involving medical imaging or not, there is no shortage of types of visual primitives that be selected to match their characteristics. Moreover, the potential applications of this approach extend beyond the visual domain, as the only prerequisite of the method is the availability of primitives that can be parameterized and ‘rendered’ in a differentiable fashion.
Acknowledgments
This research was supported by the statutory funds of Poznan University of Technology and the Polish Ministry of Science and Higher Education, grants no. 0311/SBAD/0740 and 0311/SBAD/0752.
@@ Appendix A The data
Table 3 provides the detailed information about the dataset created for this study.
Table 3: The 30 digital slides from the BRD database Bio (2023) of GTEx images used in this study. These examinations were selected using the search keyword “Thyroid gland“. Each entry contains the image in SVS format and metadata, including the case ID, sex, and age (presented below). Another field in the database contains pathology review comments, for instance, “2 pieces, no abnormalities“. This global information was used to determine the class of the sample (seen in the Class column). In total, ten images per class were selected – five for training, two for validation, and three for testing. The last column describes the number of patches extracted from a digital slide.
@@ Appendix B Configurations and hyperparameter settings
The Masked Mean Squared Error (MMSE) function described in Section 2.2 was used to not force the model to generate shapes that are only partially present in the input image.
@@
Appendix C Statistical testing of classification results
We conducted statistical tests (Fig. 9), which confirm that our encoder-modeler-renderer models are better than the autoencoders on accuracy and weighed average of F1-score.

### v3

@@ 4.4 Stage 1: Image reconstruction
Table 1 compares the performance of configurations in terms of four reconstruction metrics comparing the produced output image to the input image, all calculated on the test set: Mean Absolute Error (MAE), Mean Square Error (MSE), Structural Similarity Index Measure (SSIM) by Wang et al. (2004), and Masked MSE (MMSE), i.e. the actual reconstruction loss used in training (Sec.2).
Table 1: The performance of individual runs of three variants of ASR and the Baseline autoencoder. The model with the best MSE within a configuration is marked in bold.
Model MSE ↓ MAE ↓ SSIM ↑ MMSE ↓
Base_1 0.0342 0.1393 0.1936 0.0342
Base_2 0.0330 0.1370 0.1943 0.0330
Base_3 0.0327 0.1362 0.1938 0.0326
Base_4 0.0329 0.1367 0.1924 0.0329
Base_5 0.0326 0.1363 0.1921 0.0325
Mean 0.0331 0.1371 0.1932 0.0330
Regularized_1 0.0341 0.1415 0.1968 0.0341
Regularized_2 0.0341 0.1417 0.1976 0.0341
Regularized_3 0.0350 0.1431 0.1972 0.0350
Regularized_4 0.0342 0.1407 0.1976 0.0342
Regularized_5 0.0345 0.1412 0.1961 0.0345
Mean 0.0344 0.1416 0.1971 0.0344
Incremental_1 0.0342 0.1415 0.1956 0.0342
Incremental_2 0.0347 0.1424 0.1958 0.0347
Incremental_3 0.0349 0.1428 0.1968 0.0349
Incremental_4 0.0366 0.1466 0.1961 0.0366
Incremental_5 0.0346 0.1420 0.1960 0.0346
Mean 0.0350 0.1431 0.1961 0.0350
Baseline_1 0.0274 0.1178 0.2299 -
Baseline_2 0.0276 0.1187 0.2271 -
Baseline_3 0.0272 0.1179 0.2295 -
Baseline_4 0.0271 0.1174 0.2292 -
Baseline_5 0.0278 0.1202 0.2227 -
Mean 0.0274 0.1184 0.2277 -
@@
Figure 3: Reconstructions of sample test images produced by the Regularized_2 model.
@@ 4.5 Stage 2: Classification
The encoding parts of compared architectures produce image features that are subsequently used to train and test a classifier. For the three ASR variants, this process comprises two steps for a given bag $B$: (i) Querying the Encoder and the Modelers on each of the sixteen 256$\times$256 image patches from $B$, each resulting with vectors of latent features per scale and location in grids. (ii) Aggregation of the latent features simultaneously over space (the X and Y dimensions of the spatial latent) and over $B$, for each scale (0, 1, 2) and each latent variable independently. As there are 6 ellipse-controlling variables at each scale, there are 18 variables in total. These are then aggregated independently using mean and standard deviation, resulting in 36 features. The result of the above process is a tabular dataset of 1,463 bags (examples, rows) described by 36 attributes (columns) and partitioned into the three considered decision classes. Division into training, validation and test subsets remains the same.
@@
The induction algorithm. To demonstrate the explanatory potential of ASR, we choose decision trees as the predictive model to be trained on the tabular datasets prepared in the way described above (DecisionTreeClassifier implementation from the scikit-learn library (Pedregosa et al., 2011), an optimized version of the CART algorithm by Breiman et al. (1984)). For fairness of comparison, all induced classifiers are subject to extensive hyperparameter tuning, detailed in C.
@@
Results. We produce a decision tree for each ASR and Baseline model listed in Table 1, and asses its predictive capacity on the testing part of the tabular dataset. Table 2 presents the accuracy of classification, precision, recall and F1 score of each of those classifiers. Notably, all configurations of ASR outperform the Baseline models on all metrics, and do so by a large margin. The best Baseline model is not better than the worst ASR model on any metric. One may argue that this may be due the Baseline models having over 5 times more attributes at their disposal (200 vs. 36). However, the extensive hyperparameter tuning (C) provided fair footing for the Baseline configurations, allowing them to pick the optimal hyperparameter values and pruning intensity. Apparently, however, this was not sufficient for Baseline to perform well: the ‘anonymous’ features of its latent representation were not informative enough to provide for good predictive capacity. ASR, in contrast, managed to provide the decision trees with useful diagnostic information – even though at the time of Stage 1 training (Sec. 4.4) it was oblivious to the existence of decision classes.
@@
Table 2: Test-set predictive performance of each model. The model with the best accuracy and F1-score within a configuration is marked in bold.
Model Accuracy ↑ F1_Score ↑ Precision ↑ Recall ↑
Base_1 0.5864 0.5048 0.6245 0.5864
Base_2 0.7768 0.7735 0.8096 0.7768
Base_3 0.7112 0.7038 0.7164 0.7112
Base_4 0.5930 0.5523 0.5907 0.5930
Base_5 0.7615 0.7414 0.7957 0.7615
Regularized_1 0.6980 0.6738 0.7044 0.6980
Regularized_2 0.7133 0.7040 0.7040 0.7133
Regularized_3 0.6827 0.6655 0.6762 0.6827
Regularized_4 0.6674 0.6785 0.7079 0.6674
Regularized_5 0.7133 0.7075 0.7245 0.7133
Incremental_1 0.7090 0.6892 0.7471 0.7090
Incremental_2 0.6958 0.6998 0.7058 0.6958
Incremental_3 0.6521 0.6131 0.7061 0.6521
Incremental_4 0.7505 0.7511 0.7568 0.7505
Incremental_5 0.7396 0.7339 0.7651 0.7396
Baseline_1 0.4486 0.3572 0.2979 0.4486
Baseline_2 0.4311 0.3797 0.3945 0.4311
Baseline_3 0.2801 0.2342 0.3729 0.2801
Baseline_4 0.3129 0.2776 0.3928 0.3129
Baseline_5 0.5383 0.4486 0.4108 0.5383
@@
Statistical analysis of Table 2 (ANOVA and post-hoc Holm test) indicates significant superiority of each variant of ASR compared to the Baseline, on accuracy of classification and on the F1 score (see Sec. C details). There are no significant differences between the variants of ASR. Notably, the Base variant exhibits most variance between runs, while the Regularized one is most stable in this respect.
@@
Figure 4 presents the decision tree that achieved the best accuracy of classification across all ASR runs (Base_2, the second run of the Base configuration). Similarly to other trees induced from the ASR-generated latent representation (not shown here for brevity), this tree nicely delineates most examples of the Hashimoto decision class (H) already at the root of the tree, using the mean of the height ($h$) ellipse parameter at scale $j=1$. The remaining two decision classes (N, Normal and B, Benign) turn out to be much harder to differentiate. Overall, the Nodularity class turns out to be hardest to classify, which we attribute to the relatively large fraction of connective tissue and sparse, small blobs, which are difficult to reconstruct using only ellipses (Fig. 2). Additionally, the examinations from this decision class occasionally feature inflammation, which can be mismatched with some visual features of Hashimoto.
@@
More importantly, the obtained decision tree is notably small, comprising only 6 decision nodes and 7 leaves [^1] It is easy to notice that the pruning algorithm failed to fold the right-hand Hashimoto subtree, which would further reduce these numbers. Trees induced from attributes provided by other ASR instances were similarly compact. This corroborates the main hypothesis of this study, i.e. that the structural, compositional representation enforced by ASR correlates better with relevant visual features, and provides a more reliable basis for classification.
Interestingly, out of the five unique attributes used by decision nodes, only one (the mean height used by the root node) comes from scale $j=1$, and none come from the scale $j=2$. This signals that the features of large, coarse ellipses built at the lowest scale are most useful for this particular diagnostic scenario. To verify this conjecture, we performed feature importance analysis by assessing the gain of the Gini criterion for each attribute (a.k.a. Gini importance), and averaged it over the final, optimized and pruned decision trees obtained from all 15 instances of ASR variants from Table 1). Figure 9 shows the result of this analysis, which confirms that scale 0 features are most important, followed by scale 1 and then scale 2. In the group of color features ($a_{r}$, $a_{g}$, $a_{b}$), the standard deviation of the green color component at scale 0 turns out to be particularly informative; this may be caused by the transmissive nature of the imaging process, in which changes in the green component result in different shades of purple. Among the geometric features, the standard deviation of the orientation angle of the ellipse (d) at scale 0 proves most informative.
The transparency of ASR can be further exploited by tracing the decision-making process back to the specific instances of ellipses in the input image, i.e. those that supported the specific condition(s) on the path traversed in the decision tree.
@@ 5 Conclusions
We provided preliminary evidence for the usefulness of a new representation learning approach, in which the learner is forced to ‘explain’ the data in terms of higher-level concepts. This allows forming compact, informative and explainable representations, which are not only capable of conveying the essence of the content of the input image, but also constructing potent predictors with high explanatory power.
ASR offers natural means for providing a learning system with domain knowledge in the form of the repertoire of visual primitives. For the class of images considered in this study, elliptical shapes proved sufficient to outperform a conventional DL baseline. It is likely, however, that more sophisticated graphical representations could further improve the quality of reconstruction and the accuracy of the resulting classifiers; for instance, the Fourier-based shape representations used in (Krawiec and Nowinowski, 2024) offer more flexibility in shape modeling, and thus could likely convey more relevant information.
@@ Appendix A The data
The Biospecimen Research Database (BRD, ) (BRD) that served as the source of data in this study is a large repository of digital slides containing, among others, 893 Whole Slide Images (WSIs) of the thyroid gland, each at least 50k $\times$ 50k pixels in size and typically covering a specimen of at least 10 by 10 mm (see example slide in Fig. 5). The slides are available in Aperio SVS format, which we parsed using the slideio library (Janowczyk, 2020) that allows reading in both image data and metadata. Though dimensions of images vary significantly in the sample, their spatial resolution is the same (0.4942 $\mu$m/pix). The apparent magnification (AppMag) is also constant across the dataset (20). Figure 5 presents an example of Thyroid Gland WSI from the BRD.
Table 3 lists the examinations selected for the dataset used in the experimental part of this study. The selected sample is publicly available under the following address: https://tinyurl.com/asr-gtex.
Figure 7 presents the distribution of patients’ attributes in the selected sample, including factorization into decision classes.
Table 3: The 30 digital slides from the BRD database BRD of GTEx images used in this study. These examinations were selected using the search keyword “Thyroid gland“. Each entry contains the image in SVS format and metadata, including the case ID, sex, and age (presented below). Another field in the database contains pathology review comments, for instance, “2 pieces, no abnormalities“. This global information was used to determine the class of the sample (seen in the Class column). In total, ten images per class were selected – five for training, two for validation, and three for testing. The last column describes the number of patches extracted from a digital slide.
@@ Appendix B Configurations and hyperparameter settings
The Masked Mean Squared Error (MMSE) function described in Section 2 was used to not force the model to generate shapes that are only partially present in the input image.
@@
Figure 5: Example Thyroid Gland Whole Slide Image from the Biospecimen Research Database (BRD) BRD .
@@
Appendix C Classification (Stage 2)
We optimize the tree induction algorithm by exhaustively iterating over 30 combinations of hyperparameter values: node impurity measure (gini, entropy), maximum height of the tree (3, 4, 5, 7, $\infty$), and the minimum number of samples per leaf (5, 10, 20). For each of those combinations, the accuracy of classification is assessed using a 5-fold cross validation conducted on the training part of the tabular dataset. The combination that maximizes this measure is subsequently used to build a decision tree from the entire training set; however, to further maximize the predictive capability, this tree undergoes adaptive pruning using the minimal cost-complexity method. The intensity of pruning is controlled by the ccp_alpha parameter of DecisionTreeClassifier, with 0 being a neutral value (no pruning) and larger values causing folding of terminal subtrees into leaves. We tune ccp_alpha to maximize the classification accuracy on the validation set. The tree pruned in this way is subsequently tested on the test set.
Figure 7 presents the distribution in the tabular dataset created in Stage 2 of the experiment, to demonstrate its high similarity to the distribution of image patches presented in Fig. 6.
Figure 8 summarizes the outcomes of statisical analysis with ANOVA and post-hoc Holm test, which confirm that our encoder-modeler-renderer models are better than the autoencoders on accuracy and weighed average of F1-score.
@@ Appendix C Classification (Stage 2)
Appendix D Attribute importance
Figure 9 presents the heatmap of relative importance of attributes of visual primitives (in columns) extracted by ASR at particular spatial scales (in rows).
Figure 9: The importance of each of 36 attributes averaged over 15 decision trees obtained with ASR. The columns correspond to variables that control the rendering: the horizontal and vertical scaling ($w$ and $h$), rotation angle $d$, and RGB color components $a_{r}$, $a_{g}$, and $a_{b}$. The values of parameters from a particular scale are aggregated using mean and standard deviation (as seen in the rows).